OpenStudy (anonymous):

PLEASE HELP PLEASE what are the x-intercept(s) of the graph of y + 30= x^2 +2 ??

OpenStudy (amistre64):

xint is when y=0

OpenStudy (amistre64):

y + 30= x^2 +2 0 + 30= x^2 +2 30= x^2 +2

OpenStudy (anonymous):

And now solve for x

OpenStudy (anonymous):

How!!

OpenStudy (anonymous):

We start from here: 30=x^2+2 Now we substract 2 both sides. What do we get?

OpenStudy (anonymous):

Oh, 28= x^2

OpenStudy (anonymous):

But nothing squares to 28

OpenStudy (anonymous):

Great! And now we take the square root at both sides: \[x=\pm \sqrt{28}\]

OpenStudy (anonymous):

We can simplify a bit more using this: 28=4*7 Any idea?

OpenStudy (anonymous):

Four times seven is 28

OpenStudy (anonymous):

Factoring...? I'm confused

OpenStudy (anonymous):

@rafabc02

OpenStudy (anonymous):

Sorry, I'm back. a sec please

OpenStudy (anonymous):

\[\sqrt{28}=\sqrt{4*7}=\sqrt{4}*\sqrt{7}\] And then sqrt(4)=?

OpenStudy (anonymous):

2??

OpenStudy (anonymous):

Good! So we get this: \[x=\pm 2*\sqrt{7}\] There we have the two x intercepts: \[x=2*\sqrt{7}\] and \[x=-2*\sqrt{7}\]

OpenStudy (anonymous):

Any doubt?

OpenStudy (anonymous):

@jrosesweet Was this useful?

OpenStudy (anonymous):

@rafabc02 omg I wrote the question wrong

OpenStudy (anonymous):

It's supposed to be y+30 = x^2 +x

OpenStudy (anonymous):

Ok, no worries! We do the same: Plug x=0 and solve for y.

OpenStudy (anonymous):

No sorry Plug y=0 and solve for x

OpenStudy (anonymous):

0+30=x^2+x And now?

OpenStudy (anonymous):

30=x^2+x

OpenStudy (anonymous):

@rafabc02

OpenStudy (anonymous):

now substrac 30 both sides

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

X^2+x-30....?

OpenStudy (anonymous):

Idk

OpenStudy (anonymous):

x^2+x-30=0 Remember the quadratic formula? Take a look at the 3rd example: http://tutorial.math.lamar.edu/Classes/Alg/SolveQuadraticEqnsII.aspx

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

So put a b and c into the quadratic formula .

OpenStudy (anonymous):

That's it! And which are the values for a, b and c?

OpenStudy (anonymous):

A= 1 b=1 c=30

OpenStudy (anonymous):

@rafabc02